# A lack of LAEs within 5Mpc of a luminous quasar in an overdensity at z=6.9: potential evidence of quasar negative feedback at protocluster scales Key Words.: Cosmology: reionization – Galaxies: high-redshift – Galaxies: quasars: individual:VIK J2348–30541 Introduction

### v1

@@ 4.1 Overdensity of LAEs
Considering the number of candidates (section 3.1) and usable area of each field (section 2), we find the surface density of LAEs around VIK J2348–3054 to be deg, whilst in CDFS we only find a surface density of deg to the same depth (see §2.3 for details). If we define as a density consistent with a field density, then this quasar is sitting within an overdense region with an overdensity factor of , where the uncertainty corresponds to the 68 confidence interval. We note that given this highly skewed distribution, the probability of the field not being overdense (i.e, the probability that ) is 3. To further explore the overdensity factor and make sure the result is not driven by the CDFS density, we compare the Ly- luminosity function of our sources against that from the combined CDFS and COSMOS fields from Hu et al. (2019) and the combined LAGER 4-field luminosity function from Wold et al. (2022) in Figure 9. We define the luminosity function as:
@@
where is the effective volume of our survey– cMpc–determined by taking into account the survey area and the redshift range probed, and is the bin-width = 0.1 . We also include a detection completeness factor——as described in Hu et al. (2019) and Wold et al. (2022). This factor is calculated by inserting, and then recovering, false sources in the narrowband. The factor is shown in Figure 8 as a function of magnitude. Specifically, we use the best-fit error function shown in the Figure to estimate the luminosity function.
@@
While the overdensity is not clear in the two faintest bins, this may be due to incompleteness from our selection function. Further observations, both deeper and in other bands (as has been done for the LAGER survey, e.g., Hu et al. 2019) are likely required to determine the source of this. To estimate , we fit the amplitude of the luminosity functions from Hu et al. (2019) and Wold et al. (2022) to our data, without considering the two faintest bins. We find that the candidates around VIK J2348–3054 have a higher space density than expected from both luminosity functions; with compared to Hu et al. (2019), and compared to Wold et al. (2022). This again indicates that the quasar lives in an overdense environment.
@@
Interestingly, we are only able to find this overdensity due to the large size of our field, as there appears to be a distinct lack of sources in the vicinity of the quasar (see Figure 7). Figure 10 shows the LAE candidates’ radial distribution from the quasar. The blue shaded region represents the average CDFS surface density of deg, with an uncertainty determined by considering the Poisson single-sided, upper limit for (Gehrels, 1986). Radial bins with no detections have been given an upper limit of 1.846 objects which is again the upper limit for a single-sided Poisson distribution. Figure 10 also shows the maximum radius that objects are expected to collapse into a cluster at , 75 cMpc (9.5 pMpc) expected at this redshift (Overzier et al., 2009) and is also shown as the outer ring in Figure 7. Figure 10 shows that the field is overdense throughout all the area covered by our images, except within the inner 5pMpc from the quasar. We discuss this further in the next section.
@@ 4.2 Central suppression of Ly- emitters
Figure 7 shows the on-sky distribution of LAE candidates. The inner shaded region represents the quasar proximity zone—the theoretical region around a quasar where its UV radiation would have ionized a pure hydrogen ISM (Fan et al., 2006). Mazzucchelli et al. (2017b) determined the radius of the proximity zone of VIK J2348–3054 to be 2.64 pMpc. The Figure also shows ring with the largest radius that Overzier et al. (2009) found to contain overdensities. What is striking about the on-sky distribution is the discernible lack of candidates towards the center of the image, around the quasar position (also see Figure 10). In fact, the nearest candidate to the quasar is 0.27 away, i.e., 5.15 pMpc. This distance is indicated by the dashed ring in Figure 7. If we take the area outside of the maximum radius at which objects are expected to collapse into clusters at , i.e, the region beyond the outer ring in Figure 7, where we do not expect the quasar to have much, if any, influence, and use it to estimate the average field density, then the Poisson probability of finding the inner region void of LAEs is 1.2. However, when coupled with the fact that this underdense region is centered on the already rare luminous high-z quasar, we can conclude it is highly unlikely that this is a chance occurrence, but instead shows a real dearth of LAEs. Furthermore, we would naively expect the quasar to be close to the center of overdensity, where the environment should be more concentrated than in the outskirts, making this underdensity even more striking. In the following sections we discuss potential explanations for this observed lack of LAEs in the vicinity of VIK 2348-3054.
@@ 4.2.1 Star formation suppression due to negative feedback
The isotropic UV intensity at the Lyman limit is then and at 2 pMpc and 5 pMpc respectively, where erg cm s Hz sr and is the UV intensity at the Lyman limit, determined by dividing the local flux density by sr.
@@
Different quasar properties would influence the total observable suppression. This might explain the diversity of results within the literature, for example, a more luminous quasar would be able to ionize higher mass haloes. Another factor would be the age of the quasar and the light travel time. Quasars need time to ionize their surroundings. The longer the quasar has been turned on, the larger the suppression radius, up until . Finally, the covering factor of the quasar would directly influence how many ionizing photons could escape the AGN and affect the IGM. Therefore we would expect a relationship between covering factor and overdensity.
@@ 4.2.3 Other possible physical mechanisms
On the other hand, Willott et al. (2005), Kim et al. (2009), and Mazzucchelli et al. (2017a), raised the possibility that candidates around the quasar might be dust obscured, thereby resulting in non-detections. However, explaining our results would require a gradient in the amount of dust as a function of the distance to the quasar, which seems unlikely. Submm interferometric surveys within the inner 5 pMpc might be able to confirm or deny this effect. Although, mapping such a large area with, e.g., ALMA, would be observationally expensive. For example, ALMA observations of VIK J2348–3054 by Venemans et al. (2016), which cover an area of only 0.13 armin area, found no companion galaxies.
@@ 4.3 Comparison to other studies
Bañados et al. (2013) and Mazzucchelli et al. (2017a) searched for LAEs around ULAS J0203+0012 and PSO J215.1512–16.0417 respectively, and also performed Lyman-break selections around their respective targets. They found that the number of LBG candidates was consistent with blank fields. It is worth noting that both used the FOcal Reducer/low dispersion Spectrograph 2 (FORS2) on the Very Large Telescope (VLT) which has a FoV of 37 arcmin and they did not use an optimum filter set up, which may have influenced results.
@@ 5 Summary and Conclusion
In this paper we aimed to address important observational factors we believe have influenced previous studies which searched for overdensities around high-redshift quasars. Specifically, we searched for LAEs over an area of 3.26 deg, much larger than that probed by previous studies (see Figure 11), constituting the largest FoV searched for LAEs or LBGs around a high-redshift quasar to date. Additionally, we use a narrow-band that covers Ly- at the systemic redshift of a quasar determined from [CII] observations, to avoid biases due to uncertain redshift estimates from, e.g., CIV and MgII (see Decarli et al., 2018).
We found 38 LAE candidates around VIK J2348–3054 and applied our same selection criteria to the CDFS, images obtained with the same setup by Hu et al. (2019), for comparison. After degrading the CDFS data to match our shallower depths, we found only two LAE candidates. Adjusting for the differences in the usable areas of both fields, this implies that VIK J2348–3054 resides in an environment that is overdense by a factor of . The space density of our candidates also sit well above the luminosity functions of both Hu et al. (2019) and Wold et al. (2022) by a factor of = 112, further supporting the presence of a large overdensity.
@@ Acknowledgements.
Based on observations at Cerro Tololo Inter-American Observatory, NSF’s NOIRLab (NOIRLab Prop. ID 2021B-0905; PI: R. Assef), which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation. TSL, RJA, and AP acknowledge support from grant CONICYT + PCI + INSTITUTO MAX PLANCK DE ASTRONOMIA MPG190030. RJA acknowledge support from ANID BASAL project FB210003. RJA was supported by FONDECYT grant numbers 1191124 and 123171, and by the ANID BASAL project FB210003. JXW acknowledges support from the science research grant from the China Manned Space Project with No. CMS-CSST-2021-A05.References
@@
- Wold et al. (2019a) Wold, I. G. B., Kawinwanichakij, L., Stevans, M. L., et al. 2019a, ApJS, 240, 5
- Wold et al. (2019b) Wold, I. G. B., Kawinwanichakij, L., Stevans, M. L., et al. 2019b, ApJS, 240, 5

### v2

@@ 3.3 Star formation rates
However, it is worth noting that these star formation rate estimates are highly uncertain given that we do not know the intrinsic Ly- luminosity due to star formation and the observed Ly- luminosity is most likely a fraction of the intrinsic total.
@@ 4.1 Overdensity of LAEs
Considering the number of candidates (section 3.1) and usable area of each field (section 2), we find the surface density of LAEs around VIK J2348–3054 to be deg-2, whilst in CDFS we only find a surface density of deg-2 to the same depth (see Sect. 2.3 for details). If we define as a density consistent with a field density, then this quasar is sitting within an overdense region with an overdensity factor of , where the uncertainty corresponds to the 68 confidence interval. We note that given this highly skewed distribution, the probability of the field not being overdense (i.e, the probability that ) is 3. To further explore the overdensity factor and make sure the result is not driven by the CDFS density, we constructed a Ly- luminosity function (see the values used in Table 1) and compared it to the combined CDFS and COSMOS fields from Hu et al. (2019) and the combined LAGER 4-field luminosity function from Wold et al. (2022) in Figure 8. We define the luminosity function as:
@@
where is the effective volume of our survey– cMpc3–determined by taking into account the survey area and the redshift range probed, and is the bin-width = 0.1 . We also include a detection completeness factor——as described in Hu et al. (2019) and Wold et al. (2022). This factor is calculated by inserting, and then recovering, false sources in the narrowband. The factor is shown in Figure 8 as a function of magnitude. Specifically, we use the best-fit error function shown in the Figure to estimate the luminosity function. To further constrain which equivalent widths we are sensitive to we also investigated what range of equivalent widhts would pass our selection criteria for a given Ly- luminosity. In particular: we assumed zero flux in the i-band; for each bin in Figure 8 we calculated the percentage of synthetic LAEs that would pass our selection criteria, with equivalent widths ranging from 1 - 200 ; we then divide the luminosity function by this completeness correction factor, which results in the blue points in Figure 8. Our limit-assumption correction suggests an even larger overdensity. However, the uncommon shape of the LF observed after correction suggests that there is likely a more complex distribution of EWs, and the correct LF sits in between the black and blue dots.
@@
While the overdensity is not clear in the two faintest bins, this may be due to incompleteness from our selection function. Further observations, both deeper and in other bands (as has been done for the LAGER survey, e.g., Hu et al. 2019) are likely required to determine the source of this. To estimate , we fit the amplitude of the luminosity functions from Hu et al. (2019) and Wold et al. (2022) to our data, without considering the two faintest bins. We find that the candidates around VIK J2348–3054 have a higher space density. The luminosity functions from Hu et al. (2019) and Wold et al. (2022) are in tight agreement (as is evident in Figure 8), and both imply an overdensity of . These luminosity functions were corrected for selection incompleteness whereas our data has not been corrected in the same way. This has the consequence of the measured value representing a lower limit, which again, indicates that the quasar lives in an overdense environment.
@@
Interestingly, we are only able to find this overdensity due to the large size of our field, as there appears to be a distinct lack of sources in the vicinity of the quasar (see Figure 6). Figure 9 shows the LAE candidates’ radial distribution from the quasar. The blue shaded region represents the average CDFS surface density of deg2, with an uncertainty determined by considering the Poisson single-sided, upper limit for (Gehrels, 1986). Radial bins with no detections have been given an upper limit of 1.846 objects which is again the upper limit for a single-sided Poisson distribution. Figure 9 also shows the maximum radius that objects are expected to collapse into a cluster at , 75 cMpc (9.5 pMpc) expected at this redshift (Overzier et al., 2009) and is also shown as the outer ring in Figure 6. Figure 9 shows that the field is overdense throughout all the area covered by our images, except within the inner 5pMpc from the quasar. We discuss this further in the next section.
| Luminosity Range | Counts | ||
| [erg s | Mpc-3 | Mpc-3 | |
| 42.6 – 42.7 | 2 | -3.62 | 0.31 |
| 42.7 – 42.8 | 7 | -3.07 | 0.16 |
| 42.8 – 42.9 | 11 | -2.88 | 0.13 |
| 42.9 – 43.0 | 9 | -2.96 | 0.14 |
| 43.0 – 43.1 | 2 | -3.61 | 0.31 |
| 43.1 – 43.2 | 3 | -3.44 | 0.25 |
| 43.2 – 43.3 | 2 | -3.62 | 0.31 |
| 43.3 – 43.4 | 0 | - | - |
| 43.4 – 43.5 | 2 | -3.62 | 0.31 |
| 43.5 – 43.6 | 0 | - | - |
| 43.6 – 43.7 | 0 | - | - |
@@ 4.2 Central suppression of Ly- emitters
Figure 6 shows the on-sky distribution of LAE candidates. The inner shaded region represents the quasar proximity zone—the theoretical region around a quasar where its UV radiation would have ionized a pure hydrogen ISM (Fan et al., 2006). Mazzucchelli et al. (2017b) determined the radius of the proximity zone of VIK J2348–3054 to be 2.64 pMpc. The Figure also shows ring with the largest radius that Overzier et al. (2009) found to contain overdensities. It is worth noting that there is a small observational artifact in the form of striations, to the south of the quasar, which overlaps the proximity zone. This is caused by a relatively strong main reflection pattern in DECam, possibly by moonlight. However, the area affected is very small and this effect had no impact on LAE candidate selection.
What is striking about the on-sky distribution is the discernible lack of candidates towards the center of the image, around the quasar position (also see Figure 9). In fact, the nearest candidate to the quasar is 0.27∘ away, i.e., 5.15 pMpc. This distance is indicated by the dashed ring in Figure 6. If we take the area outside of the maximum radius at which objects are expected to collapse into clusters at , i.e, the region beyond the outer ring in Figure 6, where we do not expect the quasar to have much, if any, influence, and use it to estimate the average field density, then the Poisson probability of finding the inner region void of LAEs is 1.2. However, when coupled with the fact that this underdense region is centered on the already rare luminous high-z quasar, we can conclude it is highly unlikely that this is a chance occurrence, but instead shows a real dearth of LAEs. Furthermore, we would naively expect the quasar to be close to the center of overdensity, where the environment should be more concentrated than in the outskirts, making this underdensity even more striking. In the following sections we discuss potential explanations for this observed lack of LAEs in the vicinity of VIK 2348-3054.
@@ 4.2.1 Star formation suppression due to negative feedback
The isotropic UV intensity at the Lyman limit is then and at 2 pMpc and 5 pMpc respectively, where erg cm-2 s-1 Hz-1 sr-1 and is the UV intensity at the Lyman limit, determined by dividing the local flux density by sr.
@@
Different quasar properties would influence the total observable suppression. This might explain the diversity of results within the literature, for example, a more luminous quasar would be able to ionize higher mass haloes. Another factor would be the age of the quasar and the light travel time. Quasars need time to ionize their surroundings. The longer the quasar has been turned on, the larger the suppression radius, up until . Finally, the covering factor of the quasar ( along with the orientation of the host galaxy, or how off-centered the quasar is) would directly influence how many ionizing photons could escape the AGN and affect the IGM. Therefore we would expect a relationship between covering factor and overdensity.
@@ 4.2.2 Cosmic variance
4.2.3 Sensitivity
As mentioned, suppression of Ly- emission should occur in galaxies with halo masses less than . However, if a survey is unable to detect faint LAEs which fall into this regime, then it is possible that the faint LAEs do exist around the central quasar, but are just undetected. It is important to note that there have been confirmed detections of LAEs near quasars that LAEs near quasars have been spectroscopically confirmed using JWST Iani et al. (2024). This suggests that follow up observations at much higher sensitivities, possibly using JWST, are needed. However, if sensitivity were the reason for the lack of detected LAEs around VIKJ2348–3054 then this would then raise the question as to why we don’t detect any high-mass LAEs in the immediate vicinity of the quasar. So while sensitivity is an important component to consider, it does not fully explain our observed distribution of LAEs.
@@ 4.2.4 Other possible physical mechanisms
On the other hand, Willott et al. (2005), Kim et al. (2009), and Mazzucchelli et al. (2017a), raised the possibility that candidates around the quasar might be dust obscured, thereby resulting in non-detections. However, explaining our results would require a gradient in the amount of dust as a function of the distance to the quasar, which seems unlikely. Submm interferometric surveys within the inner 5 pMpc might be able to confirm or deny this effect. Although, mapping such a large area with, e.g., ALMA, would be observationally expensive. For example, ALMA observations of VIK J2348–3054 by Venemans et al. (2016), which cover an area of only 0.13 arcmin2 area, found no companion galaxies. Recently, Li et al. (2023) searched for submm galaxies around quasars—including VIK J2348–3054—using the Submillimeter Common-User Bolometre Array-2 (SCUBA-2) on the James Clerk Maxwell Telescope, which has a much larger 15’ field-of-view. This field of view translates to 5pMpc. Li et al. (2023) only found a single SMG around VIK J2348–3054, 0.5 pMpc away from the quasar, suggesting that dust obscuration may not be the cause of the central lack of LAEs.
@@ 4.3 Comparison to other studies
Bañados et al. (2013) and Mazzucchelli et al. (2017a) searched for LAEs around ULAS J0203+0012 and PSO J215.1512–16.0417 respectively, and also performed Lyman-break selections around their respective targets. They found that the number of LBG candidates was consistent with blank fields. It is worth noting that both used the FOcal Reducer/low dispersion Spectrograph 2 (FORS2) on the Very Large Telescope (VLT) which has a FoV of 37 arcmin2 and they did not use an optimum filter set up, which may have influenced results.
@@ 5 Summary and Conclusion
In this paper we aimed to address important observational factors we believe have influenced previous studies which searched for overdensities around high-redshift quasars. Specifically, we searched for LAEs over an area of 3.26 deg2, much larger than that probed by previous studies (see Figure 10), constituting the largest FoV searched for LAEs or LBGs around a high-redshift quasar to date. Additionally, we use a narrow-band that covers Ly- at the systemic redshift of a quasar determined from [CII] observations, to avoid biases due to uncertain redshift estimates from, e.g., CIV and MgII (see Decarli et al., 2018).
We found 38 LAE candidates around VIK J2348–3054 and applied our same selection criteria to the CDFS, images obtained with the same setup by Hu et al. (2019), for comparison. After degrading the CDFS data to match our shallower depths, we found only two LAE candidates. Adjusting for the differences in the usable areas of both fields, this implies that VIK J2348–3054 resides in an environment that is overdense by a factor of . The space density of our candidates also sit well above the luminosity functions of both Hu et al. (2019) and Wold et al. (2022) by a factor of = 40.5, further supporting the presence of a large overdensity.
@@ Acknowledgements.
We thank the anonymous referee for the valuable comments. Based on observations at Cerro Tololo Inter-American Observatory, NSF’s NOIRLab (NOIRLab Prop. ID 2021B-0905; PI: R. Assef), which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation. TSL, RJA, and AP acknowledge support from grant CONICYT + PCI + INSTITUTO MAX PLANCK DE ASTRONOMIA MPG190030. RJA acknowledge support from ANID BASAL project FB210003. RJA was supported by FONDECYT grant numbers 1191124 and 123171, and by the ANID BASAL project FB210003. CM acknowledges support from Fondecyt Iniciacion grant 11240336 and the ANID BASAL project FB210003. JXW acknowledges support from the science research grant from the China Manned Space Project with No. CMS-CSST-2021-A07.References
@@
- Wold et al. (2019) Wold, I. G. B., Kawinwanichakij, L., Stevans, M. L., et al. 2019, ApJS, 240, 5
@@
Appendix A Additional Material
| ID | R.A. | Dec. | i | z | NB964 | SFR | ||||
| (J2000.00) | (J2000.00) | (deg) | (cMpc) | (pMpc) | (ergs s-1) | ( yr-1) | ||||
| LAE-1 | 23:50:39.46 | -31:47:26.41 | 0.99 | 149.20 | 18.89 | >27.15 | 24.97 0.16 | 24.11 0.12 | 7.6 2.7 | 4.7 1.7 |
| LAE-2 | 23:49:25.39 | -31:46:34.95 | 0.89 | 133.93 | 16.95 | >27.15 | 26.07 0.43 | 24.73 0.22 | 6.6 2.7 | 4.1 1.7 |
| LAE-3 | 23:49:45.53 | -31:39:01.46 | 0.79 | 118.57 | 15.01 | >27.15 | 25.74 0.33 | 24.59 0.19 | 6.6 2.8 | 4.0 1.7 |
| LAE-4 | 23:48:22.23 | -31:37:53.92 | 0.73 | 109.47 | 13.86 | >27.15 | 25.93 0.38 | 24.46 0.17 | 9.0 2.7 | 5.5 1.7 |
| LAE-5 | 23:48:17.08 | -31:30:18.94 | 0.61 | 90.77 | 11.49 | >27.15 | 26.09 0.45 | 23.59 0.07 | 26.5 2.7 | 16.4 1.7 |
| LAE-6 | 23:46:51.65 | -31:19:29.93 | 0.56 | 83.48 | 10.57 | >27.15 | 25.85 0.36 | 24.37 0.16 | 9.8 2.8 | 6.1 1.7 |
| LAE-7 | 23:52:04.95 | -31:16:20.60 | 0.84 | 126.08 | 15.96 | >27.15 | 25.38 0.24 | 23.50 0.07 | 25.5 2.8 | 15.7 1.7 |
| LAE-8 | 23:44:51.34 | -31:14:01.78 | 0.86 | 128.78 | 16.30 | >27.15 | 25.99 0.42 | 24.08 0.12 | 15.0 2.8 | 9.2 1.7 |
| LAE-9 | 23:49:09.11 | -31:11:23.68 | 0.31 | 47.13 | 5.97 | >27.15 | 25.84 0.36 | 24.53 0.18 | 7.7 2.8 | 4.8 1.7 |
| LAE-10 | 23:48:38.18 | -31:10:25.49 | 0.27 | 40.72 | 5.15 | >27.15 | 26.12 0.47 | 24.49 0.18 | 9.4 2.8 | 5.8 1.7 |
| LAE-11 | 23:52:11.44 | -31:10:09.82 | 0.82 | 123.43 | 15.62 | >27.15 | 25.55 0.28 | 24.47 0.18 | 6.9 2.8 | 4.2 1.8 |
| LAE-12 | 23:50:09.67 | -31:00:28.43 | 0.36 | 53.98 | 6.83 | >27.15 | 25.54 0.28 | 24.20 0.14 | 10.7 2.8 | 6.6 1.7 |
| LAE-13 | 23:52:50.08 | -30:59:23.05 | 0.92 | 138.21 | 17.50 | >27.15 | 25.81 0.36 | 24.37 0.16 | 9.7 2.9 | 6.0 1.8 |
| LAE-14 | 23:46:52.33 | -30:57:07.02 | 0.36 | 54.65 | 6.92 | >27.15 | 26.07 0.44 | 24.66 0.22 | 7.3 2.8 | 4.5 1.7 |
| LAE-15 | 23:52:39.27 | -30:51:50.73 | 0.88 | 132.02 | 16.71 | >27.15 | 25.44 0.27 | 24.32 0.16 | 8.1 2.9 | 5.0 1.8 |
| LAE-16 | 23:51:42.32 | -30:50:46.97 | 0.68 | 101.71 | 12.87 | >27.15 | 25.33 0.23 | 24.46 0.16 | 5.5 2.7 | 3.4 1.7 |
| LAE-17 | 23:45:30.76 | -30:50:42.20 | 0.66 | 98.31 | 12.44 | >27.15 | 25.86 0.37 | 24.72 0.22 | 5.8 2.8 | 3.6 1.7 |
| LAE-18 | 23:45:38.56 | -30:41:06.23 | 0.66 | 99.34 | 12.57 | >27.15 | 25.69 0.32 | 24.28 0.14 | 10.4 2.7 | 6.4 1.7 |
| LAE-19 | 23:50:56.96 | -30:40:16.95 | 0.56 | 84.55 | 10.70 | >27.15 | 26.03 0.43 | 24.07 0.12 | 15.3 2.8 | 9.5 1.7 |
| LAE-20 | 23:46:20.25 | -30:40:03.94 | 0.53 | 79.67 | 10.08 | >27.15 | 25.38 0.24 | 24.57 0.19 | 4.6 2.8 | 2.8 1.7 |
| LAE-21 | 23:48:12.27 | -30:32:34.32 | 0.37 | 55.15 | 6.98 | >27.15 | 26.02 0.42 | 24.22 0.14 | 12.9 2.8 | 7.9 1.7 |
| LAE-22 | 23:44:45.07 | -30:31:07.12 | 0.90 | 135.50 | 17.15 | >27.15 | 26.05 0.45 | 24.52 0.18 | 8.8 2.8 | 5.4 1.7 |
| LAE-23 | 23:47:37.50 | -30:28:28.35 | 0.47 | 70.89 | 8.97 | >27.15 | 25.53 0.27 | 24.60 0.20 | 5.2 2.8 | 3.2 1.7 |
| LAE-24 | 23:44:31.64 | -30:27:35.79 | 0.97 | 145.88 | 18.47 | >27.15 | 25.39 0.24 | 24.41 0.17 | 6.6 2.8 | 4.1 1.8 |
| LAE-25 | 23:47:38.55 | -30:25:47.75 | 0.51 | 76.79 | 9.72 | >27.15 | 25.33 0.23 | 24.49 0.18 | 5.1 2.8 | 3.2 1.7 |
| LAE-26 | 23:44:36.91 | -30:25:04.43 | 0.98 | 146.42 | 18.53 | >27.15 | 25.95 0.40 | 23.97 0.11 | 17.0 2.8 | 10.5 1.7 |
| LAE-27 | 23:48:21.02 | -30:24:04.19 | 0.50 | 75.52 | 9.56 | >27.15 | 26.00 0.41 | 24.52 0.18 | 8.5 2.8 | 5.3 1.7 |
| LAE-28 | 23:47:21.19 | -30:22:53.54 | 0.58 | 87.27 | 11.05 | >27.15 | 25.63 0.30 | 23.84 0.10 | 18.1 2.8 | 11.2 1.7 |
| LAE-29 | 23:46:27.28 | -30:21:51.57 | 0.70 | 105.43 | 13.35 | >27.15 | 25.52 0.27 | 24.40 0.16 | 7.6 2.8 | 4.7 1.7 |
| LAE-30 | 23:49:27.05 | -30:21:25.98 | 0.58 | 86.77 | 10.98 | >27.15 | 25.65 0.30 | 24.61 0.19 | 5.9 2.7 | 3.6 1.7 |
| LAE-31 | 23:50:27.20 | -30:21:13.39 | 0.68 | 102.61 | 12.99 | >27.15 | 25.50 0.27 | 24.26 0.14 | 9.5 2.8 | 5.8 1.7 |
| LAE-32 | 23:45:02.91 | -30:12:32.87 | 1.03 | 153.76 | 19.46 | >27.15 | 25.96 0.39 | 24.73 0.21 | 6.1 2.7 | 3.8 1.7 |
| LAE-33 | 23:49:34.65 | -30:12:40.49 | 0.73 | 108.83 | 13.78 | >27.15 | 25.66 0.31 | 24.52 0.18 | 6.9 2.8 | 4.3 1.7 |
| LAE-34 | 23:50:23.24 | -30:10:45.41 | 0.82 | 123.58 | 15.64 | >27.15 | 26.13 0.48 | 24.62 0.20 | 7.9 2.8 | 4.9 1.7 |
| LAE-35 | 23:50:45.07 | -30:05:13.35 | 0.94 | 141.41 | 17.90 | >27.15 | 25.34 0.23 | 24.32 0.15 | 7.5 2.8 | 4.6 1.7 |
| LAE-36 | 23:49:43.48 | -30:03:01.26 | 0.89 | 133.31 | 16.87 | >27.15 | 25.55 0.28 | 24.31 0.15 | 9.1 2.8 | 5.6 1.7 |
| LAE-37 | 23:50:36.54 | -30:01:46.79 | 0.98 | 146.80 | 18.58 | >27.15 | 25.45 0.25 | 24.58 0.19 | 4.9 2.8 | 3.0 1.7 |
| LAE-38 | 23:48:43.58 | -29:53:13.94 | 1.02 | 152.41 | 19.29 | >27.15 | 25.80 0.36 | 24.70 0.22 | 5.7 2.9 | 3.5 1.8 |